OpenStudy (anonymous):

multiply & simplify: sqrt21 * sqrt6=

jimthompson5910 (jim_thompson5910):

sqrt(21)*sqrt(6) = sqrt(21*6) = sqrt(126) = sqrt(9*14) = sqrt(9)*sqrt(14) = 3sqrt(14) So the answer is \[\large 3\sqrt{14}\]